OpenStudy (anonymous):

Integrate e^((-3x+2sinx)/6)

OpenStudy (anonymous):

Let u = -3x+2sinx)/6 du/dx = ...

OpenStudy (anonymous):

I have written the formula but my explored break down xD First use integration by parts, integrate, derivate e^xblabla expresion and integrate 1. Then you will have another integrate in which you could do a substitution using the e exponent.

OpenStudy (anonymous):

Int ( e^u) = e^u/ du

OpenStudy (anonymous):

Still not sure about this - any help?

OpenStudy (anonymous):

correct your problem.

OpenStudy (anonymous):

\[\exp((-3x+sinx)/6)\]

OpenStudy (anonymous):

\[\exp((-3x+2sinx)/6)\]

OpenStudy (anonymous):

that what I mean, sorry

OpenStudy (anonymous):

I need to integrate it. I have made some progress but need to check it

OpenStudy (anonymous):

it has no solution i think it was e^-3x (sinx/6)

OpenStudy (anonymous):

i have got -2 e ^(sinx/6) -(x/2)

OpenStudy (anonymous):

Can anybody tell if this seems correct?

OpenStudy (anonymous):

no i certainly tell you that your problem has a mistake

OpenStudy (anonymous):

Ok. How?

OpenStudy (anonymous):

You need a range limits of integration

OpenStudy (anonymous):

sinx can not be in power of e^